OpenStudy (anonymous):

Solve the system of equations using elimination or substitution. 2x+3y=-2 5x+2y=-27

OpenStudy (anonymous):

2x+3y=-2-----------(1 5x+2y=-27----------(2 Let use the elimination method to solve the given system of equations. Multiplying eq (1by -2 and that of (2 by 3 and adding them we find -4x-6y=4-----------(1 15x+6y=-81---------(2 ------------------------ 11x =-77 i.e. x= -77/11 i.e. x=-7 Substituting x=-7 in any of the euation say eq (1 we find 2(-7)+3y=-2 i.e. -14+3y=-2 i.e. 3y = 14-2 i.e. 3y = 12 i.e. y= 12/3 i.e. y= 4 Thus {x, y} = { -7, 4} is the solution set of the given equations

OpenStudy (anonymous):

@sarcos11

OpenStudy (anonymous):

thnx